​Roswell Company has the following information on one of its vehicles purchased on January 1,

Roswell Company has the following information on one of its vehicles purchased on January 1, 

Year 1: 

Vehicle cost $50,000 

Useful life, years, estimated 5

Useful life, miles, estimated 100.000 

Salvage value, estimated $10,000 

Actual miles driven: Year 1 30,000 

Year 2 20,000 

Year 3 15,000 

Year 4 25,000 

Year 5 12.000 

No estimates were changed during the life of the asset. 


The Year 3 depreciation expense for Roswell's vehicle using the sum-of-the-years'-digits (SYD) method was 

A. $6,000 B. $8,000 C. $10,000 D. $16,000 


Using the activity method, what was Roswell's Year 5 depreciation expense? 

A. $4,000 B. $4,800 C. $5,000 D. $6.000 
Answer #1

14) Sum of the year digit = 5+4+3+2+1 = 15

Depreciation expense = (50000-10000)*3/15 = 8000

So answer is b) $8000

15) Accumulated Depreciation expense = (50000-10000)*90000/100000 = 36000

Year 5 Depreciation expense = 40000-36000 = 4000

So answer is a) $4000